OpenStudy (andriod09):

Okay people, I need help with Quadratic equations, i.e. \[ax^{2}+bx+c=0\] I have 5 and i will only put 1 here. so keep up!

OpenStudy (cgreenwade2000):

Sounds easy enough. DO you know the Quadratic formula?

OpenStudy (andriod09):

\[4x^{2}+9x+2-0\] I know that the answer is: \[x=\frac{ -9\pm \sqrt{81-4(4)(2)} }{ 8 }\]

OpenStudy (andriod09):

How do i get from point A to point B? aka from the equation to the answer

OpenStudy (andriod09):

yea, thats what it says in my book. I do the Life of Fred series and i am doing Quadratic Equation.

OpenStudy (andriod09):

rly? its that easy?!?!?!?!?! omfg!

OpenStudy (cgreenwade2000):

Take it a step further and come up with two different answers android. I believe in you.

OpenStudy (andriod09):

how are ther two different answers?

OpenStudy (andriod09):

Honestly, ill say it in the 3 different launguages i know. English:no Spanish:no ASL:shankes head

OpenStudy (andriod09):

ik but, i don't get how you got them. how do you go from\[-9\pm \frac{ \sqrt{81-32} }{ 8 } \to -9\pm \frac{\sqrt{49} }{ 2 } \]

OpenStudy (anonymous):

For \[ax^2+bx+c=0=4x^2+9x+2\] \[a=4,b=9,c=2\] using the quadratic formula \[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] substituting in \[a=4,b=9,c=2\] \[x=\frac{-9\pm \sqrt{9^2-4(4)(2)}}{2(4)}\] \[x=\frac{-9\pm \sqrt{81-32}}{8}\] \[x=\frac{-9 \pm \sqrt{49}}{8}\] \[x=\frac{-9\pm 7}{8}\] \[x=\frac{-9+7}{8}\] \[x=\frac{-2}{8}=\frac{-1}{4}\] and \[x=\frac{-9-7}{8}\] \[x=\frac{-16}{8}=-2\]

OpenStudy (andriod09):

okay, delete that long post and spearate the posts by a \[\to\] that get put into each other so its easier to read.